# UVGS: Reimagining Unstructured 3D Gaussian Splatting using UV Mapping

### v1

@@ 3.3 Mapping Networks
Our goal is to bring the extracted 3DGS attribute maps to a common feature space to better represent the object collectively and to make the 14-channel UVGS representations $U\in\mathbb{R}^{M\times N\times 14}$ work with the widely available image based foundation models. To accomplish this, we map it to a 3-channel image which can be easily processed by the existing architectures while also maintaining the spatial correspondence. We designed a simple yet effective multi-branch CNN to extract features from different attributes of the UVGS and map it to a 3-channel feature-rich image, termed Super UVGS. The structured UVGS maps provides local and global features that can be learned by a CNN.
Forward Mapping The first layer is a set of three mapping branches for position, transform, and appearance ($\phi^{f}_{P},~{}\phi^{f}_{T},~{}\phi^{f}_{A}$) respectively. We refer to them as position, transformation, and appearance branch. The position branch takes the mean position ($\sigma$) as an input and processes it to give a position feature map $M_{P}$. Similarly, the transformation branch takes the rotation ($r$) and scale ($s$) together to generate another feature map $M_{T}$. The last, appearance branch takes the color ($c$) and opacity ($o$) together to produce another feature map $M_{A}$. All the three features maps from position, transformation, and appearance branch are concatenated to get a final feature map, before passing them to the next module, called the Central branch. The central branch ($\phi^{f}_{C}$) is composed of multiple hidden Convolution layers, where each layer is followed by BatchNorm and ReLu activation. The last layer of central branch is activated using $tanh$ to ensure the Super UVGS does not take any ambiguous value resulting in gradient explosion or undesired artifacts. The obtained Super UVGS $S$ representation squeezes all the 3DGS attributes to a 3 dimensional image while also maintaining local and global structural correspondence among them.
@@
Inverse Mapping We design an inverse mapping network that aims to map the obtained 3-channel Super UVGS image $S\in\mathbb{R}^{M\times N\times 3}$ back to the UVGS maps to obtain each of the five different 3DGS attributes $\{\sigma,~{}r,~{}s,~{}o,~{}c\}$. The inverse mapping network simply follows the forward mapping network architecture in the reverse order, where at first, we put the Central branch ($\phi_{iC}$) followed by attribute specific position, transformation, and appearance branches ($\phi_{iP},~{}\phi_{iT},~{}\phi_{iA}$). We provide more details behind mapping network in the supplementary.
Branched mapping layers: The rationale behind using branched mapping layers in both forward and reverse mapping networks is to prevent the incompatibility issues arising due the the different value distribution of 3DGS attributes. Note that the disparate distributions of values within each set of attributes in 3D Gaussian Splatting, (i.e., mean position, transformation, and color), pose a challenge for the model when processed collectively. For instance, neighboring Gaussians in UVGS maps show smooth changes in position and color values but have large variations in rotation, scale, and opacity values. This results in gradient anomalies and slow training. To address this, we propose a multi-branch network architecture, where attribute-specific branches implicitly learn to process these distinct attribute specific properties, focusing on their unique features before passing them to the central branch. The central branch receives a concatenated stack of processed attributes and exploits the correlation between different attributes by extracting local feature correspondence between them. This information is then mapped to a 3-channel Super UVGS image, effectively capturing the complex relationships between the various attributes. This approach enables our network to manage diverse attribute distributions, resulting in faster convergence, improved accuracy, and specialized processing for each attribute set.
Reconstruction Losses: Since the obtained UVGS maps have both local and global features, we opted for image-based losses to train the overall architecture. We use a set of Mean Squared Error (MSE) and Learned Perceptual Image Patch Similarity (LPIPS) [63] between the obtained UVGS from spherical mapping $U$ and the predicted UVGS $\hat{U}$ from inverse mapping network. We calculate the LPIPS loss over four attributes of 3DGS including mean position ($\sigma$), view independent color ($c$), scale ($s$), and rotation ($r$). The overall LPIPS loss for UV maps can be written as a linear sum of individual attribute loss terms as:
@@
The overall loss function for the training can be written as: $\mathcal{L}_{uvgs}=\mathcal{L}_{mse}+\lambda.\mathcal{L}_{UV-lpips}$ where $\lambda$ is a scalar and varied from 00 to $10$ during the course of training.
@@ 4 Experiments
Figure 5: Figure shows a wide variety of high-quality unconditional generation result from our method using Latent Diffusion Model. We train an LDM to randomly sample Super UVGS images from random noise. The Super UVGS can be converted to 3DGS object using inverse mapping network and inverse spherical projection.
@@
Baselines & Metrics: To evaluate the quality of reconstructed 3DGS objects from both Super UVGS image and Autoencoder latent space to 3D, we use PSNR and LPIPS [62]. The aim is to convert the given 3DGS object to UVGS, and then to Super UVGS, and further to Autoencoder’s latent space and calculate the metrics from the reconstructions at every step to prove the proposed method doesn’t significantly impact the quality of reconstructions, while also providing a structurally meaningful representation that is much compact and easier to use with existing image based models. We compare the generational capabilities of our method against various conditional and unconditional SOTA 3D object generation method including the ones using multiview rendering for optimization DiffTF [3], Get3D [13], methods trying to give structural representation to Gaussians, GaussianCube [59], TriplaneGaussian [67], and general purpose SOTA large 3D content generation models like DreamGaussian [47], LGM [49], and EG3D [4]. We also compare the quality of our generation results using FID and KID.
Figure 6: Comparison of unconditional 3D asset generation on the cars category with SOTA methods. Figure shows that DiffTF [3] produces low-quality, low-resolution cars lacking detail. While Get3D [13] and GaussianCube [59] achieve higher resolution, they suffer from 3D inconsistency, numerous artifacts, and lack richness in 3D detail. In contrast, our method generates high-quality, high-resolution objects that are 3D consistent with sharp, well-defined edges.
Mapping Network Training Details We train the forward and inverse mapping networks to project the obtained UVGS maps $U\in\mathbb{R}^{M\times N\times 14}$ to a 3-channel Super UVGS image $S\in\mathbb{R}^{M\times N\times 3}$, and back to the reconstructed UVGS maps $\hat{U}\in\mathbb{R}^{M\times N\times 14}$. We provide an in-depth discussion of all implementation details in the supplementary material.
@@
[…]
3DGS $34.6~{}/~{}34.2$ ${}/~{}0.02$ 0.02 00
UVGS (@K=1) $31.3~{}/~{}31.1$ ${}/~{}0.06$ 0.06 $53.0$
UVGS (@K=2) $32.8~{}/~{}31.9$ $0.04~{}/~{}0.05$ $45.6$
UVGS (@K=4) $34.2~{}/~{}33.2$ $0.02~{}/~{}0.03$ $33.3$
Super UVGS (@K=1) $31.2~{}/~{}31.1$ $0.07~{}/~{}0.08$ $89.7$
AE (@K=1) $30.9~{}/~{}30.8$ $0.07~{}/~{}0.09$ $99.5$
VAE (@K=1) $30.6~{}/~{}30.9$ $0.07~{}/~{}0.09$ $99.5$
VQVAE (@K=1) $30.3~{}/~{}30.1$ $0.08~{}/~{}0.10$ $99.7$
@@ 4.1 UVGS AutoEncoder and 3DGS Compression
The obtained Super UVGS image is a structurally meaningful representation that can have various applications in the generation and reconstruction of new 3D assets because it contains features that can be learned by the existing image based models. Through our experiments, we show that a 3-channel Super UVGS image can be directly reconstructed using a pretrained image based Autoencoders or VAEs without any fine-tuning. We tested on three different models including image AE, KL-VAE [19], VQVAE [51] and each performed quite well without any significant quality loss. The reconstruction PSNR and LPIPS values are presented in Table 1. This means we can now leverage the powerful compression capabilities of image based Autoencoders to compress the storage requirements of 3DGS by more than $99\%$. We have shows the storage comparison results in Table 1. It is interesting to note that the Super UVGS representation itself can be used to compress the memory requirement for storing 3DGS object by up to $89.7\%$.
Figure 7: We compare the performance of our model against various SOTA methods for text-conditional object synthesis. Our method not only generated high-quality 3D assets for simpler objects, but also for complicated objects with intricate geometry.
@@ 4.2 Unconditional & Conditional Generation
We aim to show the effectiveness of the Super UVGS representation for directly generating 3DGS objects from a learned latent space. We consider Super UVGS images as a compact and structured proxy for representing 3DGS objects as it maintains the 3D object information, while also providing features for learning. The existing methods fail to directly generate a large number of Gaussians (e.g. 100k+) to represent objects with sufficient quality either due to the lack of 3D generative model architectures that support such large number of unstructured points, or due to the lack of a large 3DGS dataset [45, 49, 67, 59, 31]. We leverage the Super UVGS representation and use the existing 2D image generative models like VAEs [11] or Diffusion Models [40] for this task. Specifically, to train a generative model capable of randomly sampling new high-quality 3DGS assets for various downstream tasks, we use an unconditional Latent Diffusion Model (LDM) [40] on the obtained Super UVGS images. As illustrated in Section 4.1, we can use a pretrained image VAE for mapping the Super UVGS image to a latent space and reconstruct back. Hence, we only need to train an LDM on the latent space. More implementation details are provided in the supplementary.
@@
Unconditional LDM: To design a generative model capable of randomly sampling new high-quality 3DGS assets for various downstream tasks, we train a unconditional Latent Diffusion Model [40] on the learned Super UVGS images. Following [40, 61, 37], we use DDIM [16] for faster and consistent sampling with up to 1000 time steps used in the forward diffusion process, and 20 during denoising. Once trained, the model is used to randomly sample Super UVGS images, resulting in high quality 3DGS assets. Results are presented in Fig 5. We demonstrate that our method inherently learns to generate multiview consistent images due to the powerful Super UVGS representation unlike most prior works due to the rendering-based losses.
@@
Conditional LDM: Similar to unconditional generation, we also trained the text-conditioned LDM following the SD’s [40, 61, 37] pipeline and using the predicted text for our dataset. The trained model can be used to generate high-quality text-conditioned 3DGS assets that are multiview consistent. The results are demonstrated in Fig 6.
@@
The above experiments proves the effectiveness of our proposed Super UVGS representation in 3D object synthesis using the widely available 2D image based models. It also highlights the fact that this compact 3-channel Super UVGS representation stores not just the spatial correspondence among different pixels, but also stores rich 3D information of the objects. This way, we can easily convert a 3D asset generation problem into a 2D image generation problem without the use of any complex 3D architectures to handle large amount of unstructured and permutation invariant 3DGS points, and neither relying upon computationally expensive multiview rendering or SDS loss.
@@
Leveraging the powerful Super UVGS representation, we present in Fig. 8 one of the first experiments on inpainting 3D Gaussian Splatting directly without using any multiview rendering or distilling information from diffusion models. We try to recover the missing Gaussians by leveraging the denoising capabilities of LDM and trying to predict the missing corresponding parts of the Super UVGS image. We believe, this can have potential applications in sparse view reconstruction. More details on inpainting architecture are given in the supplementary material.
@@ 4.4 Ablation Studies & Discussion
We conduct exhaustive ablation studies to justify some of our framework’s design choices including the effect of branching in mapping networks, the use of single layer UVGS maps, and the resolution of UVGS maps. We performed our experiments on our custom Objaverse [10] 3DGS dataset and evaluate the performance of our model in terms of PSNR, SSIM, and LPIPS. The results are presented in Table 3. From the table, it can be seen that by using four layers of UVGS maps (K=4), we can almost match the reconstruction quality of fitted 3DGS results, while we realized that simply with a single layer UVGS maps, we are able to maintain the overall geometry and appearance of the object of our dataset with a PSNR of more than 30. We also compared the reconstruction performance of our method with and without using branching in the mapping network. It can be clearly seen that using branching network significantly increases the reconstruction quality from Super UVGS space. The main reasoning behind this is the specialization of attributes that the branching provides to individually process each attribute first.
@@
Get3D $68.95$ $4.38$ DreamGaussian [47] $28.51$
DiffTF $89.14$ $8.73$ Shap. E [7] $30.76$
EG3D $74.51$ $6.62$ LGM [49] $31.24$
GaussianCube $48.44$ $4.89$ GaussianCube [59] $30.27$
@@
UVGS @$K=1$ $31.1$ $0.06$ $512\times 512~{}(@K=1)$ $31.1$ $0.08$
UVGS @$K=2$ $31.9$ $0.05$ $256\times 256~{}(@K=1)$ $28.2$ $0.23$
UVGS @$K=4$ $33.2$ $0.03$ Without Branching $27.8$ $0.31$

### v4

@@ 3.3 Mapping Networks
Our goal is to bring the extracted UVGS maps to a common feature space to better represent the object collectively and to make the 14-channel UVGS representations $U\in\mathbb{R}^{M\times N\times 14}$ work with the widely available image based foundation models. To accomplish this, we map it to a 3-channel image which can be easily processed by the existing architectures while also maintaining the spatial correspondence. We design a simple yet effective multi-branch CNN to extract features from different UVGS attributes and map them to a 3-channel feature-rich image, termed Super UVGS. The structured UVGS maps provides local and global features that can be learned by a CNN.
Forward Mapping The first layer is a set of three mapping branches for position, transform, and appearance ($\phi^{f}_{P},~\phi^{f}_{T},~\phi^{f}_{A}$) respectively. We refer to them as position, transformation, and appearance branch. The position branch takes the mean position ($\sigma$) as an input and processes it to give a position feature map $M_{P}$. Similarly, the transformation branch takes the rotation ($r$) and scale ($s$) together to generate another feature map $M_{T}$. The last, appearance branch takes the color ($c$) and opacity ($o$) together to produce another feature map $M_{A}$. All the three features maps from position, transformation, and appearance branch are concatenated to get a final feature map, before passing them to the next module, called the Central Branch. The central branch ($\phi^{f}_{C}$) is composed of multiple hidden Convolution layers, where each layer is followed by BatchNorm and ReLu activation. The last layer of the central branch is activated using tanhtanh to ensure the Super UVGS does not take any ambiguous value resulting in gradient explosion or undesired artifacts. The obtained Super UVGS $S$ representation squeezes all the 3DGS attributes to a 3 dimensional image while also maintaining local and global structural correspondence among them.
@@
Inverse Mapping We design an inverse mapping network that aims to map the obtained 3-channel Super UVGS image $S\in\mathbb{R}^{M\times N\times 3}$ back to the UVGS maps to obtain each of the five different 3DGS attributes $\{\sigma,~r,~s,~o,~c\}$. The inverse mapping network simply follows the forward mapping network architecture in the reverse order, where at first, we put the Central Branch ($\phi_{iC}$) followed by attribute specific position, transformation, and appearance branches ($\phi_{iP},~\phi_{iT},~\phi_{iA}$). We provide more details on mapping network in the supplementary.
Branched mapping layers: The rationale behind using branched mapping layers in both forward and reverse mapping networks is to prevent the incompatibility issues arising due the the different value distribution of 3DGS attributes. Note that the disparate distributions of values within each set of attributes in 3D Gaussian Splatting, (i.e., mean position, transformation, and color), pose a challenge to the model when processed collectively. For instance, neighboring Gaussians in UVGS maps show smooth changes in position and color values but typically have large variations in rotation, scale, and opacity values. This results in gradient anomalies and slow convergence. To address this, we propose a multi-branch network architecture, where attribute-specific branches implicitly learn to process these distinct attribute specific properties, focusing on their unique features before passing them to the central branch. The central branch receives a concatenated stack of processed attributes and exploits the correlation between them by extracting local feature correspondences. This information is then mapped to a 3-channel Super UVGS image, effectively capturing the complex relationships between the various attributes. This approach enables our network to manage diverse attribute distributions, resulting in faster convergence, improved accuracy, and specialized processing for each attribute set.
Reconstruction Losses: Since the obtained UVGS maps have both local and global features, we opted for image-based losses to train the overall architecture. We use a set of Mean Squared Error (MSE) and Learned Perceptual Image Patch Similarity (LPIPS) [zhang2018perceptual] between the obtained UVGS from spherical mapping $U$ and the predicted UVGS $\hat{U}$ from inverse mapping network. We calculate the LPIPS loss over four attributes of 3DGS including mean position ($\sigma$), view independent color ($c$), scale ($s$), and rotation ($r$). The overall LPIPS loss for UV maps can be written as a linear sum of individual attribute loss terms as:
@@
The overall loss function for the training can be written as: $\mathcal{L}_{uvgs}=\mathcal{L}_{mse}+\lambda.\mathcal{L}_{UV-lpips}$ where $\lambda$ is a scalar and varied from 0 to $10$ during the course of training.
@@ 4 Experiments
Figure 5: Figure shows a wide variety of high-quality unconditional generation result from our method using Latent Diffusion Model. We train an LDM to randomly sample Super UVGS images from random noise. The Super UVGS can be converted to 3DGS object using inverse mapping network and inverse spherical projection. The unconditional generation model was trained on Objaverse dataset.
@@
Baselines & Metrics: To evaluate the quality of reconstructed 3DGS objects from both Super UVGS image and Autoencoder latent space to 3D, we use PSNR and LPIPS [lpips2018]. The aim is to convert the given 3DGS object to UVGS, and then to Super UVGS, and further to Autoencoder’s latent space and calculate the metrics from the reconstructions at every step to prove the proposed method doesn’t significantly affect the quality of reconstructions, while also providing a structurally meaningful representation that is much compact and easier to use with existing image based models. We compare the generational capabilities of our method against various conditional and unconditional SOTA 3D object generation method including the ones using multiview rendering for optimization DiffTF [difftf2023], Get3D [get3d2022], methods trying to give structural representation to Gaussians, GaussianCube [gaussiancube2024], and general purpose SOTA large 3D content generation models like DreamGaussian [dreamgaussian2023], LGM [lgm2025], and EG3D [eg3d2022]. We also compare the quality of our generation results using FID and KID.
Figure 6: Comparison of unconditional 3D asset generation on the cars category with SOTA methods. Figure shows that DiffTF [difftf2023] produces low-quality, low-resolution cars lacking detail. While Get3D [get3d2022] achieve higher resolution, it suffers from 3D inconsistency, numerous artifacts, and lacks 3D detail. Similar issues are found in GaussianCube [gaussiancube2024] along with symmetric inconsistency in the results. In contrast, our method generates high-quality, high-resolution objects that are 3D consistent with sharp and well-defined edges.
Mapping Network Training Details We train the forward and inverse mapping networks to project the obtained UVGS maps $U\in\mathbb{R}^{M\times N\times 14}$ to Super UVGS image $S\in\mathbb{R}^{M\times N\times 3}$, and back to the reconstructed UVGS maps $\hat{U}\in\mathbb{R}^{M\times N\times 14}$. We provide an in-depth discussion of all implementation details in the supplementary material.
@@
Method PSNR($C/F$) LPIPS($C/F$) CP($\%$)
3DGS 34.6/34.234.6~/~34.2 0.02/0.020.02~/~0.02 0
UVGS (@K=1) 31.3/31.131.3~/~31.1 0.06/0.060.06~/~0.06 $53.0$
UVGS (@K=2) 32.8/31.932.8~/~31.9 0.04/0.050.04~/~0.05 $45.6$
UVGS (@K=4) 34.2/33.234.2~/~33.2 0.02/0.030.02~/~0.03 $33.3$
Super UVGS (@K=1) 31.2/31.131.2~/~31.1 0.07/0.080.07~/~0.08 $89.7$
AE (@K=1) 30.9/30.830.9~/~30.8 0.07/0.090.07~/~0.09 $99.5$
VAE (@K=1) 30.6/30.930.6~/~30.9 0.07/0.090.07~/~0.09 $99.5$
VQVAE (@K=1) 30.3/30.130.3~/~30.1 0.08/0.100.08~/~0.10 $99.7$
@@ 4.1 UVGS AutoEncoder and 3DGS Compression
The obtained Super UVGS image is a structurally meaningful representation that can have various applications in the generation and reconstruction of new 3D assets as it contains features that can be learned by the existing image based models. Through our experiments, we show that a 3-channel Super UVGS image can be directly reconstructed using a pretrained image based Autoencoders or VAEs without any fine-tuning. We tested on three different models including image AE, KL-VAE [KLvae2013], VQVAE [vqvae2017] and each performed quite well without any significant quality loss. The reconstruction PSNR and LPIPS values are presented in Table 1. This means we can now leverage the powerful compression capabilities of image based Autoencoders to compress the storage requirements of 3DGS by more than $99\%$. We have shows the storage comparison results in Table 1. It is interesting to note that the Super UVGS representation itself can be used to compress the memory requirement for storing 3DGS object by up to $89.7\%$.
Figure 7: We compare the performance of our model against various SOTA methods for text-conditional object synthesis. Our method not only generates high-quality assets for simpler objects, but also for complicated objects with intricate geometry.
@@ 4.2 Unconditional & Conditional Generation
We aim to show the effectiveness of Super UVGS representation for directly generating 3DGS objects from a learned latent space. We consider Super UVGS images as a compact and structured proxy for representing 3DGS objects as it maintains the 3D object information while also providing learnable features. The existing methods fail to directly generate a large number of Gaussians (e.g. 100K+) to represent objects with sufficient quality either due to the lack of 3D generative model architectures that support such large number of unstructured points, or due to the lack of a large 3DGS dataset [splatterimage2024, lgm2025, triplanemeetsgs2024, gaussiancube2024, gsd2024]. We leverage the Super UVGS representation and use the existing 2D image generative models like Diffusion Models [ddim] for this task. Specifically, to train a generative model capable of randomly sampling new high-quality 3DGS assets for various downstream tasks, we use an unconditional Latent Diffusion Model (LDM) [stable_diffusion] on the obtained Super UVGS images. As illustrated in Section 4.1, we can use a pretrained image VAE to map the Super UVGS image to a latent space and reconstruct back. Hence, we only train a LDM on the latent space. More implementation details are provided in the supplementary.
@@
Unconditional LDM: To design a generative model capable of randomly sampling new high-quality 3DGS assets for various downstream tasks, we train an unconditional LDM [stable_diffusion] on the learned Super UVGS images. Following [stable_diffusion, controlnet, egosonics], we use DDIM [ddim] for faster and consistent sampling with up to 1000 time steps used in the forward diffusion process, and 20 during denoising. Once trained, the model is used to randomly sample Super UVGS images, resulting in high quality 3DGS assets through inverse mapping. Results are presented in Fig 5. We demonstrate that our method inherently learns to generate multiview consistent images due to the powerful Super UVGS representation unlike most prior works using rendering-based losses.
@@
Conditional LDM: Similar to unconditional generation, we also trained a text-conditioned LDM following the SD’s [stable_diffusion, controlnet, egosonics] pipeline and using the predicted text for our dataset. The trained model can be used to generate high-quality text-conditioned 3DGS assets that are multiview consistent. The results are demonstrated in Fig 6.
@@
The above experiments proves the effectiveness of our proposed Super UVGS representation in 3D object synthesis using widely available 2D image models. It also highlights that this compact 3-channel Super UVGS representation stores not just the spatial correspondence among different pixels, but also the rich 3D information of the objects. This way, we can easily convert a 3D asset generation problem into a 2D image generation problem without the use of any complex 3D architecture to handle large amount of unstructured and permutation invariant 3DGS primitives, and neither relying upon computationally expensive multiview rendering or SDS loss. Baseline comparison for unconditional and conditional generation is presented in Table 2.
@@
Leveraging the powerful Super UVGS representation, we present in Fig. 8 one of the first experiments on inpainting 3DGS directly without using any multiview rendering or distilling information from diffusion models. We try to recover the missing Gaussians by leveraging the denoising capabilities of LDM and trying to predict the missing corresponding parts of the Super UVGS image. We believe, this can have potential applications in sparse view reconstruction. More details are given in the supplementary.
@@ 4.4 Ablation Studies & Discussion
We conduct exhaustive ablation studies to justify some of our framework’s design choices including the effect of branching in mapping networks, the use of single layer UVGS maps, and the resolution of UVGS maps. We performed our experiments on our custom Objaverse [objaverse2023] 3DGS dataset and evaluate the performance of our model in terms of PSNR, SSIM, and LPIPS. The results are presented in Table 3. From the table, it can be seen that by using four layers of UVGS maps (K=4), we can almost match the reconstruction quality of fitted 3DGS results, while we realized that simply with a single layer UVGS, we are able to maintain the overall geometry and appearance of the object for our dataset with a PSNR of more than 30. We also compared the reconstruction performance of our method with and without using branching in the mapping network. It can be clearly seen that using branching network significantly increases the reconstruction quality from Super UVGS space. The main reasoning behind this is the specialization of attributes that the branching provides to individually process each attribute first.
@@
Get3D [get3d2022] $53.17$ $4.19$ DreamGaussian [dreamgaussian2023] $28.51$
DiffTF [difftf2023] $84.57$ $8.73$ Shap. E [shapeditor2024] $30.53$
EG3D [eg3d2022] $74.51$ $6.62$ LGM [lgm2025] $30.74$
GaussianCube $34.67$ $3.72$ GaussianCube [gaussiancube2024] $30.34$
@@
UVGS @K=1@K=1 $31.1$ $0.06$ $512\times 512~(@K=1)$ $31.1$ $0.08$
UVGS @K=2@K=2 $31.9$ $0.05$ $256\times 256~(@K=1)$ $28.2$ $0.23$
UVGS @K=4@K=4 $33.2$ $0.03$ Without Branching $27.8$ $0.31$